answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of rolling the sum of 6 with 2 dice?

there is a 1 to 5 probablilty that you will roll the sum of 6 with two dice.

What is the probability that the number 23 will be randomly picked out of 100?

Assuming then that there are 100 numbers, 1-100, the probability of the number 23 being randomly picked out of 100 is: 1/100 or 0.01.

If you roll one die with 6 sides what is the probability that you will roll a 3 or a 4?

The answer is 1/3. There are six possible outcomes (1 to 6) of which two (3 or 4) are favourable so the probability is 2/6 or 1/3. In general, if A and B are two events, then Pr (A or B) = Pr(A) + Pr(B0 - Pr(A and B) [the last bit is because you are double counting those events] Here Pr(A) = Pr(3) = 1/6, Pr(B) = Pr(4) = 1/6 and Pr(A and B) = Pr(3 and 4 - simultaneously) = 0 So Pr(3 or 4) = 1/6 + 1/6 + 0 = 1/3

What is sample space for tossing 2 coins?

The sample space for tossing 2 coins is (H = Heads & T = Tails): HH, HT, TH, TT

If two dice are thrown what is the probability of the sum of the numbers being a seven or an eleven?

To roll a 7, it doesn't matter what your first die rolls at, as every number on the die has a complement that will add up to 7. This means that your odds on rolling 7 are actually your odds on rolling the appropriate number on the second die. ie. 1/6. No matter what your first die is, the second can potentially give you a number that brings the total to 7.

To roll an 11 however, your odds are not as high. This is because you must roll a five and a six to hit 11. Your first die can be either of the two, but your second die must be it's complement. In this case then, your odds are 2/6 * 1/6, or 1/18.

What is probability of having 0ne boy out of 3 children?

We can write the sample space for 3 children as follows (b=boy, g=girl). bbb gbb bbg gbg bgb ggg bgg ggb The sample space is 8; each has a probability of 1/8. From above, you can see 3 instances that have 1 boy; bgg, gbg & ggb. Therefore, add 1/8 + 1/8 + 1/8 = 3/8 probability one boy having three children.

How does a discrete probability distribution differ from a continuous probability distribution?

A discrete probability distribution is defined over a set value (such as a value of 1 or 2 or 3, etc). A continuous probability distribution is defined over an infinite number of points (such as all values between 1 and 3, inclusive).

What is the probability of getting a number greater than 2 when rolling a single die once?

There are 6 possible outcome when rolling a die. Of those, 3,4,5, and 6 are greater than 2. That means 4/6 odd of getting a number greater than 2. But 4/6 is the same as 2/3 or 66.67% Also, some dice are different from the standard ones. Here is another answer someone had: 

well it depends........if it's a 1-5 die then there's only 3/5 chances which is 60% chance so you are higher likely to get more than 2 if its an 1-7 die then its about a percent that i don't know but the more-number of sides it has the more chance you get more than 2....hope you win at whatever you win or this helped on what ever you are doing

An experiment consists of flipping a coin and then rolling a die What is the probability of observing a head on the coin or a 3 on the die?

The probability of flipping a Head is the same as that for a Tail and is 1/2 or 50%. The probability of rolling a particular number on a die is 1/6 since there are 6 numbers. Combining these two probabilities (by multiplication) we have, as the combined probability 1/2 x 1/6 = 1/12 = 0.0833333333333333(the 3 recurs) which as a percentage is 8.33333333333%

You flipped a coin two times. What is the probability of getting at least one head?

With two flips of a coin you can get two heads, two tails, a head and a tail, or a tail and a head. There are a total of four different possible outcomes, and three of them have at least one head. That's 3 out of 4, or 3/4ths. It's also 0.75 which is the probability of getting at least one head with two flips of a coin. Note that as we use the term probability here, it is zero (no chance it can happen at all), or one (it must happen), or something in between. A probability appears in the form of a fraction or decimal, and has no units attached to it.

Students are required to answer 2 True of False questions and 1 multiple choice questions with 4 responses If the answers are all guesses what is the probability of getting all 3 questions correct?

The probability of correct true & false question is 1/2 and the probability correct multiple choice (four answer) question is 1/4. We want the probability of correct, correct, and correct. Therefore the probability all 3 questions correct is 1/2 * 1/2 * 1/4 = 1/16.

There are 2 White 4 Black and 6 Yellow balls in an urn If you select 2 balls what is the probability that they are the same color?

To find the solution, describe the events that will allow the same color draws which are WW or BB or YY. Since replacement is not stated, without replacement is assumed. So, we need probability of WW, BB, and YY (added together) which are (2/12*1/11) + (4/12*3/11) + (6/12*5/11) or (2/132 + 12/132 + 30/132) or 44/132 or1/3.

You flip a coin nine times and obtain ttttttttt if you flip the coin one more time what is the probability of getting h?

The probability of a flipped coin landing heads or tails will always be 50% either way, no matter how many times you flip it.

What percent of the scores will fall between 1.22 and 8.78 Mean 5 Standard Deviation 1.41?

Convert to Standard Normal Distribution the values of 1.22 & 8.78. Z = (8.78-5)/1.41 = 2.68; and Z = (1.22-5)/1.41 = -2.68. Find the area between 2.68 & -2.68 from Table. Area @ 2.68 = .9963; Area @ -2.68 = .0037. Take the difference in the areas and that is the solution. 0.9963 - 0.0037 = .9926 or 99.26% of the scores fall between 1.22 & 8.78.

The probability that an individual is left handed is 0.1 In a class of 30 students what is the probability of finding at least 5 left hander's?

This is a binomial probability distribution. The number of trials, n, equals 30; and the probability of success is p, which is 0.1. In this problem, you want the probability of at least 5, which is the complement of at most 4. We use the complement because we can subtract from 1 that probability and we will have the solution. The related link has the binomial probability distribution table which is cumulative. Per the table, at n=30, p=0.1 and x = 4; the probability is 0.825. Therefore the probability of at least 5 is 1 - 0.825 or 0.175.

If a box contains 4 red marbles seven white and 5 blue and two marbles are drawn one at a time with replacement What is the probability that both marbles are white?

Since the box contains 16 marbles, seven of them white, then the probability of drawing one white marble is 7/16. If you replace the marble and draw again, the probability of drawing another white marble is still 7/16. The net probability of drawing two white marbles, while replacing the first, is 49/256.

In craps if you play both the 6 and 8 what are your odds against rolling a 7?

In craps, the chance for rolling a 7 never changes. First roll, tenth roll - doesn't matter. The chance is always 6 out of 36 (or roughly 17%). The chance of rolling a 6 is always 5 out of 36 (or 20%). The chance of rolling an 8 is always 5 out of 36 (or 20%). The chance of hitting either a 6 or an 8 before hitting a 7 would be 10 to 6 (with the favor going towards hitting either the 6 or 8).

It is required to seat 5 men and 4 women in a row so that the women occupy the even places how many such arrangements are possible?

There are two ways to look at this problem; gender or person.

Gender (M=Male, F=Female); There is only one way and that is MFMFMFMFM.

By person, you have 9 places to fill. The first is 5 males could be placed there, the next 4 females could be placed there, the next 4 males can be placed there, the next 3 females can be placed there and so forth. So in this instance there are 5*4*4*3*3*2*2*1*1 = 2880 arrangements.

Another way is

just 5!*4!=2880.

Here 5! is for men and 4! for women.

ie. Men can be arranged in 5 ways and women can be arranged in 4 ways.

What is non-probability sampling?

Non probability sampling is where the samples are not selected randomly.

If you toss three coins 240 times how many times can you expect the coins to have three tails showing?

If we toss three coins 240 times, how many times can we expect the coins to have three tails showing?

Why the probability is have a specified chance to be selected and why the non probability does not have an equal of being selected?

Probability is the number of occurrences of a specific event divided by the total number of possible events. Suppose you took a multiple choice test with 4 possible answer's A, B, C, D. Only 1 of the 4 choices is correct; so the probability of correct is 1/4 and the probability of incorrect is 3/4. Most of the time the two probabilities will not be the same.

What is the probability of a coin tossed four times to get four heads?

The probability of a fair coin to land head is 1/2. Since for 4 flips it must land heads each time, the probability of 4 heads is 1/2 * 1/2 * 1/2 * 1/2 = 1/16.